a frog lies at the bottom of a well whose depth is 30m. It starts to climb, rising up 2/5 every day and falling back 1/5 every night. How long will it take him to climb out of the well?
1
Expert's answer
2012-10-26T11:14:16-0400
As the frog rises 2/5 every day and falling back 1/5 every night, then its 24-hour rising is

2/5 - 1/5 = 1/5 m.

So, it will take

N = 30/ (1/5) = 30*5 = 150 days

to climb out.